# Signature matrices of membranes

### v1

@@ 2 Iterated integrals
    For membranes, that is $p=2$ parameters and standard basis $e_{1},\dots,e_{d}$ of $V=\mathbb{R}^{d}$, we obtain the $\mathsf{id}$-signature according to [14]. Its level $k$ component can be stored in the $d\times\dots\times d$ array,
    $\sigma^{(k)}_{i_{1},\dots,i_{k}}$ $:=\sigma(e_{i_{1}}\otimes\dots\otimes e_{i_{k}})$
    $=\int\limits_{\begin{subarray}{c}0<s_{1}<\dots<s_{k}<1\\ 0<t_{1}<\dots<t_{k}<1\end{subarray}}\partial_{12}X_{i_{1}}(s_{1},t_{1})\dots\partial_{12}X_{i_{k}}(s_{k},t_{k})\,\mathrm{d}s_{1}\mathrm{d}t_{1}\dots\mathrm{d}s_{k}\mathrm{d}t_{k}.$
    In particular, for level $k=2$, we obtain our signature matrix $\sigma=\sigma^{(2)}$ from the introduction, (2).
  4. 4.
    When ‘replacing’ integrals by sums, and $\partial_{12}$ by its discrete counterpart, we obtain the two-parameter sums signature [18], evaluated at all diagonal matrix compositions. For this compare also [14, Example 4.16].
@@
It is clear from the definition that the signatures of any two Lipschitz continuous maps $X,Y:[0,1]^{p}\to V$ agree if $\partial_{1...p}(\alpha\circ(X-Y))=0$ for all $\alpha\in V^{*}$. For $p=1$ this is just the translation invariance of the signature, for membranes, i.e. $p=2$, this is invariance under adding terms only depending on one of the variables. In particular, if $X:[0,1]^{2}\to V$ is a membrane, we can always find a membrane $X_{\mathrm{red}}$ which vanishes on $[0,1]\times\{0\}\cup\{0\}\times[0,1]$ and has the same signature as $X$, by letting $X_{\mathrm{red}}(s,t):=X(s,t)-X(0,t)-X(s,0)+X(0,0)$. See also Lemma 3.18 and Figure 4 for an illustration of piecewise bilinear membranes.
We provide some first examples for membranes ($p=2$).
@@
    For fixed $u\in\mathbb{R}^{d}$, the $k$-th signature tensor of the $d$-dimensional bilinear membrane $[0,1]^{2}\rightarrow\mathbb{R}^{d},(s,t)\mapsto ust$ is given by
    $\sigma^{(k)}_{i_{1},\dots,i_{k}}=\frac{u_{i_{1}}\dots u_{i_{k}}}{(k!)^{2}}$ (13)
@@
    We recall the $2$-dimensional polynomial membrane $X=A\,\mathsf{Mom}^{2,2}$ from the introduction, (7). With integration over the differentials of the two-parameter moment coordinates,
@@
    In Example 3.7 and 3.11 we provide closed formulas for its level $2$ and $3$ signature tensors, respectively.
@@
Let $V,W$ be finite dimensional $\mathbb{R}$-vector spaces, and $X:[0,1]^{p}\to V$ be a Lipschitz continuous map. If $A:V\to W$ is linear, then $\sigma(AX)=\sigma(X)\circ T(A^{*})$.
@@
In particular, Lemma 2.6 implies that $\sigma^{(k)}(AX)=A^{\otimes k}(\sigma^{(k)}(X))$ for all $k$. If $V=\mathbb{R}^{d}$ we can write this in the Tucker format $\sigma^{(k)}(AX)=[\![\sigma^{(k)}(X);A,\dots,A]\!]$. Explicitly, if $k=2$ we have
$\sigma^{(k)}(AX)=A\sigma^{(k)}(X)A^{\top},$
as was already mentioned in the introduction (6). The equivariance of signatures is a key property used in [41], allowing the computation of signature tensors for various families of paths from a small collection of dictionaries, also called core tensors. In the following section we introduce such dictionaries for membranes.
@@ 3 Product membranes
    The moment membrane of multidegree $(m,n)$ in the introduction (4) is the product
@@
    where the moment path $\mathsf{Mom}^{m}$ is according to Example 2.1. Note that we sometimes omit the last vectorization, i.e., we identity $\mathbb{R}^{m\times n}$ with $\mathbb{R}^{mn}$ whenever the context allows this. For this compare also Corollary 3.6. Note that every $d$-dimensional polynomial membrane $X$ with multidegree $(m,n)$ without terms of degree $\leq(1,1)$ is of the form $X=A\circ\mathsf{Mom}^{m,n}$, where $A$ is a linear map, e.g. our running Example 2.5 with $m=n=d=2$.
@@ 3.1 Factorizations into path signatures
$\sigma(X),\alpha_{1}\otimes\dots\otimes\alpha_{k}\right>$ $\sigma(Y),\beta_{1}\otimes\dots\otimes\beta_{k}\right>$
@@
Note that the tensor product of linear forms $\sigma(X):T(V^{*})\to\mathbb{R}$, $\sigma(Y):T(V^{*})\to\mathbb{R}$ is not the tensor product in the tensor series space $T(V^{*})$; the formula above should not be confused with Chen’s formula for the signature of a concatenation of paths.
@@
If $X:[0,1]^{2}\to V\otimes W$ is a product membrane, it can be convenient to use a vectorization $\nu:V\otimes W\cong\mathbb{R}^{\dim V}$, for example to obtain an actual signature matrix. Using Lemma 2.6, we see that $\sigma(\nu X)$ then factors as
$T(\mathbb{R}^{d})\to T((V\otimes W)^{*})\to T(V^{*})\otimes T(W^{*})\to\mathbb{R}$
@@
Let $X:[0,1]\to\mathbb{R}^{m}$ and $Y:[0,1]\to\mathbb{R}^{n}$ be two paths and let $\mathsf{kron}(X,Y):[0,1]^{2}\to\mathbb{R}^{mn}$ denote the membrane that maps $(s,t)$ to the Kronecker product of the column vectors $X(s)$ and $Y(t)$. Then
$\sigma^{(2)}(\mathsf{kron}(X,Y))$ $=\sigma^{(2)}(X)\otimes\sigma^{(2)}(Y)$
@@
$C^{(2)}:=\sigma^{(2)}(\mathsf{vec}\circ\mathsf{Mom}^{2,2})=\sigma^{(2)}(\mathsf{Mom}^{2})\otimes\sigma^{(2)}(\mathsf{Mom}^{2})=\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}\otimes\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}$
as it is presented in the introduction, (5). With $S^{(2)}:=\sigma^{(2)}(X)=AC^{(2)}A^{\top}$ we provide the three missing homogeneous entries,
$S^{(2)}_{2,1}$ $=\frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{6}a_{1,1}a_{2,2}+\frac{1}{6}a_{1,1}a_{2,3}+\frac{1}{9}a_{1,1}a_{2,4}+\frac{1}{3}a_{2,1}a_{1,2}+\frac{1}{3}a_{2,1}a_{1,3}$
@@
$S^{(2)}_{1,2}$ $=\frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{3}a_{1,1}a_{2,2}+\frac{1}{3}a_{1,1}a_{2,3}+\frac{4}{9}a_{1,1}a_{2,4}+\frac{1}{6}a_{2,1}a_{1,2}+\frac{1}{6}a_{2,1}a_{1,3}$
@@
$S^{(2)}_{2,2}$ $=\frac{1}{4}a_{2,1}^{2}+\frac{1}{2}a_{2,1}a_{2,2}+\frac{1}{2}a_{2,1}a_{2,3}+\frac{5}{9}a_{2,1}a_{2,4}+\frac{1}{4}a_{2,2}^{2}+\frac{4}{9}a_{2,2}a_{2,3}+\frac{1}{2}a_{2,2}a_{2,4}$
@@
using equivariance, Lemma 2.6. The polynomial $S_{1,1}^{(2)}$ is already provided in the introduction.
We illustrate the membrane $X$ in the first row of Figure 3, for $A=\begin{pmatrix}1&-1&1&1\\ 1&1&0&-1\end{pmatrix}$. In particular, we provide the graphs of the coordinate functions $\mathsf{graph}(X_{1})$ and $\mathsf{graph}(X_{2})$, and a plot of the image $\mathsf{Im}(X)$ in $\mathbb{R}^{2}$. Here, the colors correspond to the image of rectangles in a uniform partition of $[0,1]^{2}$, e.g. $X([0,\frac{1}{3}]\times[0,\frac{1}{2}])$ in pink, $X([\frac{1}{3},\frac{2}{3}]\times[0,\frac{1}{2}])$ in orange, or $X([\frac{2}{3},1]\times[\frac{1}{2},1])$ in violet. Note that $X$ maps three sides of the square $[0,1]^{2}$ to the origin. The right-most picture shows an ”untangled” version of the membrane $X$ which has the same signature, cf. Remark 2.4.
@@
Choosing the vectorization $\nu:\mathbb{R}^{m}\otimes\mathbb{R}^{n}\to\mathbb{R}^{mn},e_{i}\otimes e_{j}\mapsto e_{n(i-1)+j}$, we see that $\mathsf{kron}=\nu\circ X\boxtimes Y$. Then we have
@@
As another consequence of Proposition 3.3 and equivariance we obtain the formula from [14, Example 4.2] for the Hadamard product of two paths:
@@
If $V=W=\mathbb{R}^{d}$ then
@@
as morphisms $T(\mathbb{R}^{d})\to\mathbb{R}$, and where $\mu:\mathbb{R}^{d}\otimes\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the componentwise multiplication map.
@@
Remark 3.9.
Proposition 3.3 and its proof immediately generalize to higher $p$-parameter patches $[0,1]^{p}\rightarrow\mathbb{R}^{d}$.
@@
With Proposition 3.3, our identification in 3.10, and the closed formula for signatures of polynomial paths in Example 2.1, we obtain a closed formula for the level $k$ tensor of the moment membrane $\mathsf{Mom}^{m,n}$ from (14),
$\langle\sigma(\mathsf{Mom}^{m,n}),(i_{1},j_{1})\dots(i_{k},j_{k})\rangle=\frac{i_{2}\dots i_{k}}{(i_{1}+i_{2})\dots(i_{1}+\dots+i_{k})}\frac{j_{2}\dots j_{k}}{(j_{1}+j_{2})\dots(j_{1}+\dots+j_{k})}$
where $i_{1},\dots,i_{k}\in[m]$ and $j_{1},\dots,j_{k}\in[n]$. For instance for the $2$-dimensional polynomial membrane from our running example Example 2.5, we obtain the level $k=3$ core tensor
$C^{(3)}:=\scriptsize\left(\begin{array}[]{cccc|cccc|cccc|cccc}\frac{1}{36}&\frac{1}{36}&\frac{1}{36}&\frac{1}{36}&\frac{1}{24}&\frac{2}{45}&\frac{1}{24}&\frac{2}{45}&\frac{1}{24}&\frac{1}{24}&\frac{2}{45}&\frac{2}{45}&\frac{1}{16}&\frac{1}{15}&\frac{1}{15}&\frac{16}{225}\\[4.0pt] \frac{1}{72}&\frac{1}{60}&\frac{1}{72}&\frac{1}{60}&\frac{1}{45}&\frac{1}{36}&\frac{1}{45}&\frac{1}{36}&\frac{1}{48}&\frac{1}{40}&\frac{1}{45}&\frac{2}{75}&\frac{1}{30}&\frac{1}{24}&\frac{8}{225}&\frac{2}{45}\\[4.0pt] \frac{1}{72}&\frac{1}{72}&\frac{1}{60}&\frac{1}{60}&\frac{1}{48}&\frac{1}{45}&\frac{1}{40}&\frac{2}{75}&\frac{1}{45}&\frac{1}{45}&\frac{1}{36}&\frac{1}{36}&\frac{1}{30}&\frac{8}{225}&\frac{1}{24}&\frac{2}{45}\\[4.0pt] \frac{1}{144}&\frac{1}{120}&\frac{1}{120}&\frac{1}{100}&\frac{1}{90}&\frac{1}{72}&\frac{1}{75}&\frac{1}{60}&\frac{1}{90}&\frac{1}{75}&\frac{1}{72}&\frac{1}{60}&\frac{4}{225}&\frac{1}{45}&\frac{1}{45}&\frac{1}{36}\end{array}\right)$
of shape $4\times 4\times 4$, and with equivariance (Lemma 2.6), the level $k=3$ signature tensor $\sigma^{(3)}(X)=[\![C^{(3)};A,A,A]\!]$ of shape $2\times 2\times 2$. We provide its first homogeneous entry,
@@
Inspired by [18, Theorem 4.5], we obtain the following computational implication.
@@
For every $d$-dimensional polynomial membrane $X$ of multi-degree $(m,n)$ we can compute $\sigma^{(k)}(X)$ in $\mathcal{O}(m^{k}n^{k}+kd^{k+1}mn)$ elementary operations.
Proof.
We use the closed formula in Example 3.11 and $\mathsf{vec}^{\otimes k}$ to obtain the core tensor $C^{(k)}$ in $\mathcal{O}(m^{k}n^{k})$. The remaining follows with equivariance. ∎

### v2

@@ 2 Iterated integrals
    When ‘replacing’ integrals by sums, and $\partial_{12}$ by its discrete counterpart, we obtain the two-parameter sums signature [DS22], evaluated at all diagonal matrix compositions. For this compare also [diehl2024signature, Example 4.16].
@@
It is clear from the definition that the signatures of any two membranes $X,Y:[0,1]^{2}\to V$ agree if $\partial_{12}(\alpha\circ(X-Y))=0$ for all $\alpha\in V^{*}$. In particular, if $X:[0,1]^{2}\to V$ is a membrane, we can always find a membrane $X_{\mathrm{red}}$ which vanishes on $[0,1]\times\{0\}\cup\{0\}\times[0,1]$ and has the same signature as $X$, by letting $X_{\mathrm{red}}(s,t):=X(s,t)-X(0,t)-X(s,0)+X(0,0)$. See also Lemma 3.17 and Figure 4 for an illustration of piecewise bilinear membranes.
We provide some first examples for membranes.
@@
    For fixed $u\in\mathbb{R}^{d}$, the $k$-th signature tensor of the $d$-dimensional bilinear membrane $X:[0,1]^{2}\rightarrow\mathbb{R}^{d},(s,t)\mapsto ust$ is given by
    $\sigma^{(k)}_{i_{1},\dots,i_{k}}(X)=\frac{u_{i_{1}}\dots u_{i_{k}}}{(k!)^{2}}$ (11)
@@
    We recall the $2$-dimensional polynomial membrane $X=A\,\mathsf{Mom}^{2,2}$ from the introduction, (5).
    With integration over the differentials of the two-parameter moment coordinates,
@@
    In Example 3.7 and 3.10 we provide closed formulas for its level $2$ and $3$ signature tensors, respectively.
@@
Let $V,\ W$ be finite dimensional $\mathbb{R}$-vector spaces, $X:[0,1]^{2}\to V$ a membrane and $A:V\to W$ a linear map. Then $\sigma(AX)=\sigma(X)\circ T(A^{*})$ where $T(A^{*})$ denotes the map $T(W^{*})\to T(V^{*})$ induced by diagonal action of $A^{*}$ on tensors.
@@
In particular, Lemma 2.6 implies that $\sigma^{(k)}(AX)=A^{\otimes k}(\sigma^{(k)}(X))$ for all $k$. If $V=\mathbb{R}^{d}$ we can write this in the Tucker format $\sigma^{(k)}(AX)=[\![\sigma^{(k)}(X);A,\dots,A]\!]$. Explicitly, if $k=2$ and $S(X)$ denotes the signature matrix, we have
$S(AX)=AS(X)A^{\top},$
as was already mentioned in the introduction (4). The equivariance of signatures is a key property used in [bib:PSS2019], allowing the computation of signature tensors for various families of paths from a small collection of core tensors. These are given by signature tensors of special paths called dictionaries. In the following section we derive dictionaries for membranes from such dictionaries for paths.
@@ 3 Product membranes
    The moment membrane of order $(m,n)$ in the introduction (2) is the product
@@
    where the moment path $\mathsf{Mom}^{m}$ is according to Example 2.1. Note that we sometimes omit the last vectorization, i.e., we identify $\mathbb{R}^{m\times n}$ with $\mathbb{R}^{mn}$ whenever the context allows this. For this compare also Corollary 3.6. Note that every $d$-dimensional polynomial membrane $X$ of order $(m,n)$ and with terms of order $\geq(1,1)$ is of the form $X=A\circ\mathsf{Mom}^{m,n}$, where $A$ is a linear map, e.g. our running Example 2.5 with $m=n=d=2$.
@@ 3.1 Factorizations into path signatures
$\sigma(X),\alpha_{1}\otimes\dots\otimes\alpha_{k}\right>$ $\sigma(Y),\beta_{1}\otimes\dots\otimes\beta_{k}\right>.$
@@
Note that the tensor product of linear forms $\sigma(X):T(V^{*})\to\mathbb{R}$, $\sigma(Y):T(V^{*})\to\mathbb{R}$ is not the tensor product in the tensor series space $T(V^{*})^{*}$; the formula above should not be confused with Chen’s formula for the signature of a concatenation of paths.
@@
If $X:[0,1]^{2}\to V\otimes W$ is a product membrane, it can be convenient to use a vectorization $\nu:V\otimes W\cong\mathbb{R}^{\dim(V)\cdot\dim(W)}$, for example to obtain an actual signature matrix. Using Lemma 2.6, we see that $\sigma(\nu X)$ then factors as
$T(\mathbb{R}^{\dim(V)\cdot\dim(W)})\to T((V\otimes W)^{*})\to T(V^{*})\otimes T(W^{*})\to\mathbb{R}$
@@
Let $X:[0,1]\to\mathbb{R}^{m}$ and $Y:[0,1]\to\mathbb{R}^{n}$ be two paths and let $\mathsf{kron}(X,Y):[0,1]^{2}\to\mathbb{R}^{mn}$ denote the membrane that maps $(s,t)$ to the Kronecker product of the column vectors $X(s)$ and $Y(t)$. Then we have the following identity between signature matrices:
$S(\mathsf{kron}(X,Y))$ $=S(X)\otimes S(Y)$
@@
$C:=\sigma^{(2)}(\mathsf{Mom}^{2,2})=\sigma^{(2)}(\mathsf{Mom}^{2})\otimes\sigma^{(2)}(\mathsf{Mom}^{2})=\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}\otimes\begin{pmatrix}\frac{1}{2}&\frac{2}{3}\\[4.0pt] \frac{1}{3}&\frac{1}{2}\end{pmatrix}$
as it is presented in the introduction, (3). With $S(X)=\sigma^{(2)}(X)=ACA^{\top}$ we provide the three missing homogeneous entries,
$S(X)_{2,1}$ $=\frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{6}a_{1,1}a_{2,2}+\frac{1}{6}a_{1,1}a_{2,3}+\frac{1}{9}a_{1,1}a_{2,4}+\frac{1}{3}a_{2,1}a_{1,2}+\frac{1}{3}a_{2,1}a_{1,3}$
@@
$S(X)_{1,2}$ $=\frac{1}{4}a_{1,1}a_{2,1}+\frac{1}{3}a_{1,1}a_{2,2}+\frac{1}{3}a_{1,1}a_{2,3}+\frac{4}{9}a_{1,1}a_{2,4}+\frac{1}{6}a_{2,1}a_{1,2}+\frac{1}{6}a_{2,1}a_{1,3}$
@@
$S(X)_{2,2}$ $=\frac{1}{4}a_{2,1}^{2}+\frac{1}{2}a_{2,1}a_{2,2}+\frac{1}{2}a_{2,1}a_{2,3}+\frac{5}{9}a_{2,1}a_{2,4}+\frac{1}{4}a_{2,2}^{2}+\frac{4}{9}a_{2,2}a_{2,3}+\frac{1}{2}a_{2,2}a_{2,4}$
@@
using equivariance, Lemma 2.6. The polynomial $S(X)_{1,1}$ is already provided in the introduction.
We illustrate the membrane $X$ in the first row of Figure 3, for the explicit choice $A=\begin{pmatrix}1&-1&1&1\\ 1&1&0&-1\end{pmatrix}$. In particular, we provide the graphs of the coordinate functions $\mathsf{graph}(X_{1})$ and $\mathsf{graph}(X_{2})$, and a plot of the image $\mathsf{Im}(X)$ in $\mathbb{R}^{2}$. Here, the colors correspond to the image of rectangles in a uniform partition of $[0,1]^{2}$, e.g. $X([0,\frac{1}{3}]\times[0,\frac{1}{2}])$ in pink, $X([\frac{1}{3},\frac{2}{3}]\times[0,\frac{1}{2}])$ in orange, or $X([\frac{2}{3},1]\times[\frac{1}{2},1])$ in violet. Note that $X$ maps three sides of the square $[0,1]^{2}$ to the origin. The right-most picture shows an “untangled” version of the membrane $X$ which has the same signature, cf. Remark 2.4.
@@
Choosing the vectorization $\nu:\mathbb{R}^{m}\otimes\mathbb{R}^{n}\to\mathbb{R}^{mn},e_{i}\otimes e_{j}\mapsto e_{n(i-1)+j}$, we see that $\mathsf{kron}(X,Y)=\nu\circ X\boxtimes Y$. Then we have
@@
As another consequence of Proposition 3.3 and equivariance we obtain the formula from [diehl2024signature, Example 4.2] for the Hadamard product of two paths:
@@
Let $\mu:\mathbb{R}^{d}\otimes\mathbb{R}^{d}\to\mathbb{R}^{d}$ denote the componentwise multiplication map. If $V=W=\mathbb{R}^{d}$ then
@@
as morphisms $T(\mathbb{R}^{d})\to\mathbb{R}$. In particular, $S(X\odot Y)=S(X)\odot S(Y)$ where $\odot$ denotes the Hadamard product of vectors resp. matrices.
@@
With Proposition 3.3, our identification in 3.9, and the closed formula for signatures of polynomial paths in Example 2.1, we obtain a closed formula for the level $k$ tensor of the moment membrane $\mathsf{Mom}^{m,n}$ from (12),
$\langle\sigma(\mathsf{Mom}^{m,n})$ $,(i_{1},j_{1})\dots(i_{k},j_{k})\rangle$
$=\frac{i_{2}\dots i_{k}}{(i_{1}+i_{2})\dots(i_{1}+\dots+i_{k})}\frac{j_{2}\dots j_{k}}{(j_{1}+j_{2})\dots(j_{1}+\dots+j_{k})}$
where $i_{1},\dots,i_{k}\in[m]$ and $j_{1},\dots,j_{k}\in[n]$. For instance for the $2$-dimensional polynomial membrane from our running example Example 2.5, we obtain the level $k=3$ core tensor, here denoted by $C^{(3)}$,
$\scriptsize\left(\begin{array}[]{cccc|cccc|cccc|cccc}\frac{1}{36}&\frac{1}{36}&\frac{1}{36}&\frac{1}{36}&\frac{1}{24}&\frac{2}{45}&\frac{1}{24}&\frac{2}{45}&\frac{1}{24}&\frac{1}{24}&\frac{2}{45}&\frac{2}{45}&\frac{1}{16}&\frac{1}{15}&\frac{1}{15}&\frac{16}{225}\\[4.0pt] \frac{1}{72}&\frac{1}{60}&\frac{1}{72}&\frac{1}{60}&\frac{1}{45}&\frac{1}{36}&\frac{1}{45}&\frac{1}{36}&\frac{1}{48}&\frac{1}{40}&\frac{1}{45}&\frac{2}{75}&\frac{1}{30}&\frac{1}{24}&\frac{8}{225}&\frac{2}{45}\\[4.0pt] \frac{1}{72}&\frac{1}{72}&\frac{1}{60}&\frac{1}{60}&\frac{1}{48}&\frac{1}{45}&\frac{1}{40}&\frac{2}{75}&\frac{1}{45}&\frac{1}{45}&\frac{1}{36}&\frac{1}{36}&\frac{1}{30}&\frac{8}{225}&\frac{1}{24}&\frac{2}{45}\\[4.0pt] \frac{1}{144}&\frac{1}{120}&\frac{1}{120}&\frac{1}{100}&\frac{1}{90}&\frac{1}{72}&\frac{1}{75}&\frac{1}{60}&\frac{1}{90}&\frac{1}{75}&\frac{1}{72}&\frac{1}{60}&\frac{4}{225}&\frac{1}{45}&\frac{1}{45}&\frac{1}{36}\end{array}\right)$
which is of shape $4\times 4\times 4$. With equivariance (Lemma 2.6), this results in the level $k=3$ signature tensor $\sigma^{(3)}(X)=[\![C^{(3)};A,A,A]\!]$ of shape $2\times 2\times 2$. We provide its first homogeneous entry,
@@
For every $d$-dimensional polynomial membrane $X$ of order $(m,n)$ we can compute $\sigma^{(k)}(X)$ in $\mathcal{O}(m^{k}n^{k}+kd^{k+1}mn)$ elementary operations.